# Zero-order Control Barrier Functions for Sampled-Data Systems with State and Input Dependent Safety Constraints

### v1

@@ V-A Constraints with high-relative degrees
Choosing the extended class $\mathcal{K}$ function $\gamma(s)=\gamma_{c}s$ for $s\in\mathbb{R}$ with $10<\gamma_{c}<$ 1, we derive the ZOCBF condition in (4) on the sampled input $u_{k}$ at sampling time $t_{k}$:
@@
Thus, we obtain a ZOCBF constraint that is linear in $u_{k}$.
Now consider another safety constraint $h_{2}(x)=10-p^{2}$. Again, $h_{2}$ is a high-order constraint. Under the same setting as above, we derive the ZOCBF condition for constraint $h_{2}$ on the sampled input $u_{k}$ at the sampling time $t_{k}$:
@@
which is a convex quadratic constraint in $u_{k}$.
The initial state for the simulated data shown in Fig.  1 is chosen to yield a safety violation under ${u_{nom}}\equiv 0$. The figure demonstrates that our method ensures safety by filtering out the unsafe nominal controller input through the ZOCBF-based safety filter.
@@ V-B Constraints involving state and input
Our second numerical example looks at rollover prevention for differential drive robots that navigate over uneven terrains. The safety condition is based on the zero-tilting moment point (ZMP), as described in [30]. This constraint depends on both the system’s state and input. To handle this constraint, we employ our proposed ZOCBF approach.
In order to develop a nonlinear dynamic model of the vehicle as it navigates 3D terrains, we derive the kinematic equations for 3D motion in an inertial frame $\mathcal{I}$. Vehicle orientation is described by a 3D rotation between the frame $\mathcal{I}$ and a body fixed frame $\mathcal{B}$. We assume that the robot maintains ground contact, which implies that the robot’s roll angle $\alpha$ and pitch angle $\beta$ equal the ground’s slope angles and ${\alpha,\beta\in(-\pi/2,\pi/2)}$. Because the terrain varies, these angles are functions of the vehicle’s positional states, denoted as $\alpha(x)$ and $\beta(x)$. The vehicle heading angle with respect to the frame ${\mathcal{I}}$ is denoted by ${{\theta}\!\in\![0,2\pi)}$.
@@
where ${[{x}^{\mathcal{I}}~{}{y}^{\mathcal{I}}]^{\top}\!\in\!\mathbb{R}^{2}}$ is vehicle position, $\omega$ represents the robot’s yaw rate, and $v$ denotes its linear velocity in frame $\mathcal{B}$. When ${\alpha=\beta=0}$, we get the unicycle model [31].
@@
where ${K_{v},\!~{}K_{\omega}\!\geq\!0}$ are the controller gains, ${[{x}_{g}~{}y_{g}]^{\top}\!\in\!\mathbb{R}^{2}}$ is the goal position, and ${d_{g}\!\triangleq\!\cos\theta\left({x}_{g}\!-\!x^{\mathcal{I}}\right)\!+\!\sin\theta\left(y_{g}\!-\!y^{\mathcal{I}}\right)}$, ${\bar{d}_{g}\!\triangleq\!-\sin\theta\left(x_{g}\!-\!x^{\mathcal{I}}\right)\!+\!\cos\theta\left(y_{g}\!-\!y^{\mathcal{I}}\right)}$. We set ${\omega=0}$ when ${x_{g}\!=\!x^{\mathcal{I}},~{}y_{g}\!=\!y^{\mathcal{I}}}$. This will be our nominal control input that is to be filtered by our ZOCBF safety filter in (10).
ZMP serves as a metric to assess the risk of tipover in robotics systems. It represents the point on the ground where the combined forces of gravity and inertia generate a moment about the plane’s normal direction, resulting in a tipping moment. Essentially, when the ZMP remains within the robot’s support polygon—the convex hull formed by its wheel or foot contacts—the system maintains stability without any tendency to tipover. When the ZMP lies outside of this convex hull, the motion is considered unsafe. And, rollover refers to the tipping of a robotic system around its lateral (roll) axis. The ZMP-based rollover constraint from [30] can be expressed as
@@
where $\mathrm{g}$ is the gravitational constant, $h_{cg}$ is the height of the robot’s center of mass, and $b$ is the robot width.
We simulated robot operation on an undulating surface, and selected a reference path that induces rollover under the control law $\mathbf{k_{d}}$ (see Fig. 3). We computed the system flow using a fourth-order RK numerical integration technique. We then solved the online optimization problem (10) using a sequential least squares programming (SLSQP) solver. The simulated closed-loop system behavior can be seen in Fig. 2 and Fig. 3. These figures show that the proposed method ensures the robot’s rollover safety by filtering out the unsafe nominal controller through the ZOCBF-based framework.
Figure 2: Simulation of a differential drive vehicle model (23) navigating on an uneven terrain, comparing the safety-filtered trajectory using the proposed ZOCBF filter (blue) with the unsafe nominal controller (green) while following a reference trajectory (black). The initial position is denoted as $x_{0}$, while the final position is denoted as $x_{f}$. The red areas along the blue path indicate unsafe routes on which the vehicle rollovers when under the nominal (unfiltered) controller. The vehicle is able to drive on rough terrain safely with the ZOCBF, whereas it is unsafe under the nominal controller.
@@ VI Conclusion
This paper introduced zero-order control barrier functions (ZOCBFs) notion for ensuring safety of sampled-data systems. Our ZOCBF approach extends conventional CBFs by effectively and straightforwardly handling safety constraints with high-relative degrees and those dependent on both states and inputs. We presented three numerical implementation methods—dynamic linearization, numerical integration, and parallel simulation—each suited to different system dynamics and computational capabilities. Through simulations involving collision avoidance and rollover prevention on uneven terrains, we demonstrated the practical effectiveness and versatility of ZOCBFs in maintaining safety-critical operations. This advancement simplifies the design and deployment of safety filters in sampled-data systems. Future work will explore ZOCBFs implementations in higher-dimensional and more complex robotic systems.

### v2

@@ V-A Constraints with high-relative degrees
Choosing the extended class $\mathcal{K}$ function $\gamma(s)=\gamma_{c}s$ for $s\in\mathbb{R}$ with $10<\gamma_{c}<$ 1, we derive the ZOCBF condition in (4) on the sampled input $u_{k}$ at sampling time $t_{k}$
@@
Thus, we obtain a ZOCBF constraint that is linear in $u_{k}$. Now consider another safety constraint $h_{2}(x)=10-p^{2}$. Again, $h_{2}$ is a high-order constraint. Under the same setting as above, we derive the ZOCBF condition for constraint $h_{2}$ on the sampled input $u_{k}$ at the sampling time $t_{k}$
@@
which is a convex quadratic constraint in $u_{k}$. The initial state for the simulated data shown in Fig.  1 is chosen to yield a safety violation under ${u_{nom}}\equiv 0$. The figure demonstrates that the ZOCBF-based safety filter keeps the system safe.
@@ V-B Constraints involving state and input
Our second example looks at rollover prevention for differential drive robots that navigate over uneven terrains. The safety condition is based on the zero-tilting moment point (ZMP), as described in [30]. This constraint depends on both the system’s state and input. To handle this constraint, we employ our proposed ZOCBF approach.
In order to develop a nonlinear dynamic model of the vehicle as it navigates 3D terrains, we derive the kinematic equations for 3D motion in an inertial frame $\mathcal{I}$. We assume that the robot maintains ground contact, which implies that the robot’s roll angle $\alpha$ and pitch angle $\beta$ equal the ground’s slope angles and ${\alpha,\beta\in(-\pi/2,\pi/2)}$. Because the terrain varies, these angles are functions of the vehicle’s positional states, denoted as $\alpha(x)$ and $\beta(x)$. The vehicle heading angle with respect to the inertia frame ${\mathcal{I}}$ is denoted by ${{\theta}\!\in\![0,2\pi)}$.
Figure 2: Simulation of a differential drive vehicle model (22) navigating on uneven terrain, comparing the safety-filtered trajectory using the proposed ZOCBF filter (blue) with the unsafe nominal controller (green) while following a reference trajectory (black). The initial position is denoted as $x_{0}$, while the final position is denoted as $x_{f}$. The red areas along the blue path indicate unsafe routes on which the vehicle rollovers when under the nominal (unfiltered) controller. The vehicle is able to drive on rough terrain safely with the ZOCBF, whereas it is unsafe under the nominal controller.
@@
where ${[{x}^{\mathcal{I}}~{}{y}^{\mathcal{I}}]^{\top}\!\in\!\mathbb{R}^{2}}$ is vehicle position, $\omega$ represents the robot’s yaw rate, and $v$ denotes linear velocity in the body frame $\mathcal{B}$. When ${\alpha=\beta=0}$, we get the unicycle model [31].
@@
where ${K_{v},\!~{}K_{\omega}\!\geq\!0}$ are the controller gains, ${[{x}_{g}~{}y_{g}]^{\top}\!\in\!\mathbb{R}^{2}}$ is the goal position, and ${d_{g}\!\triangleq\!\cos\theta\left({x}_{g}\!-\!x^{\mathcal{I}}\right)\!+\!\sin\theta\left(y_{g}\!-\!y^{\mathcal{I}}\right)}$, ${\bar{d}_{g}\!\triangleq\!-\sin\theta\left(x_{g}\!-\!x^{\mathcal{I}}\right)\!+\!\cos\theta\left(y_{g}\!-\!y^{\mathcal{I}}\right)}$. We set ${\omega=0}$ when ${x_{g}\!=\!x^{\mathcal{I}},~{}y_{g}\!=\!y^{\mathcal{I}}}$. This acts as our nominal control input to our ZOCBF safety filter in (10).
ZMP serves as a metric to assess the risk of tipover in robotics systems. It represents the point on the ground where the combined forces of gravity and inertia generate a moment about the plane’s normal direction, resulting in a tipping moment. Essentially, when the ZMP remains within the robot’s support polygon—the convex hull formed by its wheel contacts—the system does not have any tendency to tipover. When the ZMP lies outside of this convex hull, the motion is considered unsafe. Rollover refers to the tipping of a robotic system around its lateral (roll) axis. The ZMP-based rollover constraint from [30] can be expressed as
@@
where $\mathrm{g}$ is the gravitational constant, $h_{cg}$ the height of the robot’s center of mass, and $b$ the robot’s width.
We simulated robot operation on an undulating surface, and selected a reference path that induces rollover under the control law $\mathbf{k_{d}}$ (see Fig. 3). We computed the system flow using a fourth-order RK numerical integration technique. We then solved the online optimization problem (10) using a sequential least squares programming (SLSQP) solver. The simulation was run on a laptop with a 14-core Intel i7-12700H CPU and 32 GB of RAM, and completed in 4 seconds. The simulated closed-loop system behavior can be seen in Fig. 2 and Fig. 3. These figures show that the proposed method ensures the robot’s rollover safety by filtering out the unsafe nominal controller through the ZOCBF-based framework. A parallel simulation-based ZOCBF safety filter is also implemented, which shows a similar performance and is neglected due to space limit. We refer interested readers to the code webpage1 for details.
@@ VI Conclusion
This paper introduced the notion of zero-order control barrier functions (ZOCBFs) for ensuring the safety of sampled-data systems. Our ZOCBF approach extends conventional CBFs by effectively and straightforwardly handling safety constraints with high-relative degrees and those dependent on both states and inputs. We presented three numerical implementation methods—dynamic linearization, numerical integration, and parallel simulation—each suited to different system dynamics and computational capabilities. Through simulations involving collision avoidance and rollover prevention on uneven terrain, we demonstrated the practical effectiveness and versatility of ZOCBFs in maintaining safety-critical operations. This advancement simplifies the design and deployment of safety filters in sampled-data systems. Future work will explore ZOCBFs implementations in higher-dimensional and more complex robotic systems.